I am trying to help my son with this question but do not know much pre-algebra. I came up with $5.50 (your per hour) and $5.00 (friend's per hour) but do not know how to write the equation or solve for the variable(s)
Please help!!
The question is:
Write an equation to represent the following problem and solve for the variable(s):
You and your friend each work 5 hours and together earn $52.50. You are paid $0.50 per hour more than your friend. How much do you earn per hour? How much does your friend earn per hour?

Answer by rapaljer(4551)

Let x = your friend's rate per hour for 5 hours.
x+.50 = your rate per hour for 5 hours.

The total amount earned is the sum of your earnings and your friends earnings and that equals 52.50. You each worked 5 hours, so here is the equation:

5(x) + 5(x+.50) = 52.50
5x + 5x + 2.50 = 52.50
10x + 2.50 - 2.50 = 52.50-2.50
10x = 50
x= $5 per hour (your friend's rate)
x+ .50 = $5.50 per hour (your rate)

R^2 at SCC